# Multimode Entangled Squeezed Light Generation and Propagation in a Coupled-Cavity Photonic Crystal

### v2

@@ III.4 Correlation Variance in in the output CROWs
In the previous two subsections, we modeled the evolution of the pump light as it coupled into the RS and computed the density operator of the 2MSTS generated in the RS by the pump light. In this section, we model the free quantum evolution of the 2MSTS in the full system in the presence of loss but without the pump and compute the time-dependent expectation value of the CV between output CROW cavities. We accomplish this by first solving the adjoint master equation for the time-evolution of the full-system QM ladder operators, and then use the initial state of the system at time $\tilde{t}=\tilde{t}_{0}$ to calculate time dependence of the photon number and correlation variance.
@@
Consider now the CV between two cavities, one each in the signal and idler CROW $n$ lattice sites away from their respective coupling cavity. We shall denote the $n^{th}$ cavity in the signal and idler CROW as $S_{n}$ and $I_{n}$ respectively (see Fig. 1), and the CV between them as $\Delta(S_{n},\tilde{t}_{S};I_{n},\tilde{t}_{I})$, where we allow the measurement times $\tilde{t}_{S}$ and $\tilde{t}_{I}$ to be different. This CV may be expressed in terms of the quadrature operators in the two cavities as
$\Delta(S_{n},\tilde{t}_{S};I_{n},\tilde{t}_{I})$ $=\expectationvalue{\quantity[X_{S_{n}}(\tilde{t}_{S})-X_{I_{n}}(\tilde{t}_{I})]^{2}}$
$+\expectationvalue{\quantity[Y_{S_{n}}(\tilde{t}_{S})+Y_{I_{n}}(\tilde{t}_{I})]^{2}},$ (49)
@@
The task now is to determine the time evolution of the Heisenberg operators in Eqs. (50) and (51). In the previous subsection, we modeled the dynamics of the density operator for the light in the RS only. We first determine the evolution of the full-system QM ladder operators and then we express the single-defect QM ladder operators (the $a_{q}(t)$-operators in Eqs. (50, 51)) in terms of the full-system QM ladder operators.
Due to the non-orthogonality of the basis of full-system QMs the non-Hermitian free Hamiltonian describing light in the full system takes the form [28]
where $N=184$ is the total number of full-system QMs which is equal to the number of cavities, and the full-system QM ladder operators obey the following relations:
where $\stackrel{{\scriptstyle\leftrightarrow}}{{P}}$ is the inverse of $\stackrel{{\scriptstyle\leftrightarrow}}{{O}}$ of Eq. (12). The full-system QM ladder operators obey the commutation relations:
$[c_{\mu},\ b_{\mu}^{\dagger}]=\delta_{\mu\nu},$ (55)
$[b_{\mu},\ c_{\nu}^{\dagger}]=-\delta_{\mu\nu},$ (56)
$[b_{\mu},\ b_{\nu}^{\dagger}]=O_{\mu\nu},$ (57)
$[b_{\mu}^{\dagger},\ b_{\nu}^{\dagger}]=[c_{\mu},\ c_{\nu}]=0.$ (58)
Inverting Eq. (59) yields the inverse relations
Below, we shall write the time-dependent photon number and correlation variance in terms of the $a$-operators. First, though, we must determine the time-evolution of the $c$-operators, since, as we shall see, it is these (and not the $b$-operators) that display simple single-frequency time-dependence. We begin with the adjoint master equation, which describes the time-evolution of a Heisenberg picture operator $A_{H}$ in terms of the free Hamiltonian in Eq. (52) [30]:
@@
where $c_{\nu}\equiv c_{\nu}(\tilde{t}_{0})$, displaying the simple behavior we expected. For $c^{\dagger}(\tilde{t})$ we simply take the Hermitian conjugate of Eq. (62). More generally, it is easy to show that for any normally-ordered set of $c-$operators, the evolution of each operator is simply given by the result shown in Eq. (62). The time-evolution of the $b_{\nu}$’s is found via construction from the $c_{\nu}$’s to be [28]
$b_{\mu}(\tilde{t})\equiv\sum_{\nu=1}^{N}O_{\nu\mu}c_{\nu}(\tilde{t}),$ (63)
where again $b^{\dagger}(\tilde{t})$ is obtained via the Hermitian conjugate of Eq. (63). The time-evolution of the single-cavity QM operators (the $a$-operators), obtained using Eqs. (60), (62), and (63), is found to be
$\vec{a}(\tilde{t})$ $=\stackrel{{\scriptstyle\leftrightarrow}}{{W}}^{\dagger}\stackrel{{\scriptstyle\leftrightarrow}}{{O}}\stackrel{{\scriptstyle\leftrightarrow}}{{F}}(\tilde{t})\stackrel{{\scriptstyle\leftrightarrow}}{{P}}\stackrel{{\scriptstyle\leftrightarrow}}{{V}}^{\dagger}\vec{a}$ (64)
$=\stackrel{{\scriptstyle\leftrightarrow}}{{\Phi}}(\tilde{t})\vec{a},$ (65)
where $\vec{a}\equiv\quantity(a_{1},a_{2},...,a_{p})^{T}$ and $\stackrel{{\scriptstyle\leftrightarrow}}{{F}}(\tilde{t})=\text{diag}\quantity(e^{-i\tilde{\omega}_{1}(\tilde{t}-\tilde{t}_{0})},e^{-i\tilde{\omega}_{2}(\tilde{t}-\tilde{t}_{0})},...,e^{-i\tilde{\omega}_{\nu}(\tilde{t}-\tilde{t}_{0})})$. Inserting Eq. (65) into Eqs. (50, 51) and substituting the result into Eq. (49) yields, after some algebra
$\Delta^{2}(S_{n},\tilde{t}_{S};I_{n},\tilde{t}_{I})$ $=\frac{1}{2}(B_{S_{n}S_{n}}+B_{I_{n}I_{n}})+\sum_{m,m^{\prime}=1}^{N}\quantity[\phi_{S_{nm}}^{*}(\tilde{t}_{S})\phi_{S_{nm^{\prime}}}(\tilde{t}_{S})+\phi_{I_{nm}}^{*}(\tilde{t}_{I})\phi_{I_{nm^{\prime}}}(\tilde{t}_{I})]\expectationvalue{a^{\dagger}_{m}a_{m^{\prime}}}$
[…]
In Eq. (66), all the time-dependence is contained within the $\phi_{pq}(\tilde{t})$ elements, and the only quantities still unknown are the $\expectationvalue{a^{\dagger}_{p}a_{q}}$ and $\expectationvalue{a_{p}a_{q}}$ ensemble averages, which we shall now compute. These ensemble averages are evaluated at the time $\tilde{t}=\tilde{t}_{0}$ (see Fig. 7). We assume that at this time, the only light in the system is the generated light which exists exclusively in the signal and idler QMs of the RS. This enables us to write:
where $p$ and $q$ are now cavities in the RS, the $d_{j}$ ($d_{j}^{\dagger}$) are the annihilation (creation) operators for the $j^{\text{th}}$ RS QM, and the $\sigma_{pj}$ are the matrix elements of the matrix $\stackrel{{\scriptstyle\leftrightarrow}}{{\Sigma}}=\stackrel{{\scriptstyle\leftrightarrow}}{{U}}^{-1}$ (see Eq. (15). In Eqs. (67-68), $j$ and $k$ run through only the signal ($j=S$) and idler ($k=I$) modes of the RS.
We now directly evaluate the ensemble averages in Eqs. (67-68) for use in the next two sections. Using the density operator from Eq. (34), the ensemble average for a general operator $A$, acting on a two-mode squeezed state can be expressed as:
@@
Using the standard relations [39], we obtain:
@@
Thus, from Eq. (69) and Eqs. (70 - 73), we can evaluate Eqs. (67, 68). After some algebra, we obtain:
$\expectationvalue{a_{p}a_{q}}$ $=-\quantity({\sigma}^{*}_{pS}{\sigma}^{*}_{qI}+{\sigma}^{*}_{pI}{\sigma}^{*}_{qS})\quantity(n^{th}_{S}+n^{th}_{I}+1)e^{i\theta}\cosh r,\sinh r,$ (74)
$\expectationvalue{a_{p}^{\dagger}a_{q}}$ $=\sigma_{pS}{\sigma}^{*}_{qS}\quantity(n^{th}_{S}\cosh^{2}r+(1+n^{th}_{I})\sinh^{2}r)+\sigma_{pI}{\sigma}^{*}_{qI}\quantity(n^{th}_{I}\cosh^{2}r+(1+n^{th}_{S})\sinh^{2}r).$ (75)
@@
Finally, the combination of Eqs. (74, 75) with Eq. (66) yields an expression for the correlation variance between two cavities in the signal and idler output CROWs. The preceding analysis also yields an expression for the average photon number in the output CROW cavities: since Eq. (75) is normally ordered we may simply use Eq. (65) to write
$\expectationvalue{a_{n}^{\dagger}a_{n}}(t)=\sum_{q\in RS}\phi_{q,n}^{*}(t)\phi_{q,n}(t)\expectationvalue{a_{q}^{\dagger}a_{q}}.$ (76)
In the next section, we present and discuss the solutions to Eqs. (66) and (76).
IV Results
In this section we present the results obtained for the PC CCS in Fig. 1 using our theoretical approach. In particular, we present the time evolution of the average photon number and correlation variance in cavities in the three CROWs. We begin by examining the time-dependent average photon number in cavities in the three CROWs to ascertain the spatio-temporal dynamics of the generated light. Next we compute the time-dependent correlation variance between light in the signal and idler CROWs by solving Eq. (66). We will explore a variety of pumping configurations and time delays to demonstrate the utility of our approach and the performance of our coupled cavity system.
IV.1 Photon Number
In Fig. 8, we plot the photon number in select cavities in the three CROWs as a function of time. In the signal and idler CROWS, the photon number in a given cavity exhibits a series of peaks that decrease in amplitude with time. The time that the first peak arrives in each cavity increases, and its amplitude decreases as the distance of the cavity from the RS increases. This is simply due to the effects of the group velocity and scattering loss, respectively. One might expect that the period between successive peaks in the CROWS to be given by the beat period between the signal and idler modes in the RS. This is approximately the case in the pump CROW. However, complications arising from the reflections off of the CROWS and the coupling cavities results in the oscillation period in the signal and idler CROWS being closer to twice the beat period.
Another feature seen in Fig. 8, which is pertinent to the following analysis of the CROW CV, is the difference in the propagation speeds in the signal and idler CROWs. Each of the CROWs is designed to carry light at its band center. However, due to the difference in inter-cavity coupling in each of the CROWs, the group velocity at band center is different in each of the CROWs. We infer this from Fig. 2, where the bandwidths of the three CROWs are different, and so the group velocities at band center are different. Indeed, we see that the average photon number peaks earlier in the first idler cavity than the first signal cavity, though by cavity 13 the signal light begins arriving first. This phenomenon is clearly seen in Fig. 9, where we plot the photon number in three different cavities in the signal and idler CROWS as a function of time.
From Fig. 8, we also see that for cavities sufficiently far from the RS, most of the generated light is in the signal and idler waveguides, as desired. However, at early times, a substantial fraction of the generated light is in the pump CROW cavities close to the RS. Most of this light cannot propagate down the pump CROW because its frequency is outside of the pass-band of the pump CROW (see Fig. 2). Thus, it evanescently decays and its amplitude is very small beyond the third cavity in the pump CROW. The relatively large number of generated photons in the cavities of the pump CROW that are close to the RS arises due to the strong coupling of the pump CROW to the RS. We have deliberately chosen this coupling to be strong to reduce the residence time of the pump light in the RS; in studying this system and others similar to it we have found that a shorter residence time limits the generation of thermal noise from scattering loss during the nonlinear process in the RS.
Finally, we observe from Fig. 8 that that there are more photons in the signal CROW than in the idler CROW. There are two different contributors to this imbalance. The first is the result of the different coupling strengths between the RS and the signal and idler CROWs. Since the intrinsic Qs of the signal and idler RS QMs were different, the coupling between these QMs and their respective CROWs had to be tuned separately so as to achieve the same loaded $Q$ for the signal and idler RS modes. This was accomplished via the coupling cavities indicated in Fig. 1. Manipulating the loaded $Q$s of the RS QMs in this manner resulted in a difference in the out-coupling rate between the signal and idler modes. Ideally the loaded $Q$s and the out-coupling efficiency of the signal and idler RS QMs would be approximately equal, but it was not possible to achieve this for this structure, while also keeping the equal frequency spacing between the modes. The second contributor to the imbalance in the signal and idler photon number is the effect of the strongly-coupled pump CROW, which results in a more complicated time evolution of the generated light after the pump has left. Indeed, when we perform the same analysis but consider a system with the pump CROW removed, the time-dependent photon numbers in the signal and idler CROWs are nearly equal.
IV.2 Correlation Variance
[…]
In Fig. 11, we plot the global minima and maxima of the CV envelopes from the first 50 output CROW cavity pairs with $\tilde{t}_{S}=\tilde{t}_{I}$. The minima of the CV remain below unity in all cases, demonstrating entanglement far from the RS. To highlight the importance of taking into account the QM non-orthogonality in complicated PC CCSs (see Fig. (5)), we have repeated the calculation ignoring the non-orthogonality, i.e., setting $\stackrel{{\scriptstyle\leftrightarrow}}{{O}}=\stackrel{{\scriptstyle\leftrightarrow}}{{I}}$ in Eq. (65), and present the results for the CV envelope as the dashed lines in Fig. 11. Note that for cavities close to the RS, the difference between the CV minimum and unity can change by up to 16%, indicating that it is important to include the nonorthogonality of the modes in the calculations for this system.
Figure 11: Minimum and maximum of the CV envelope for pairs of output CROW cavities as a function of distance the coupling cavity (CC). Dashed lines show the results obtained assuming that the system QMs are orthogonal.
We now consider the effects of varying the pump pulse energy and width in $k$-space on the CV and the optimal measurement times. Although the naïve choice of $\tilde{t}_{S}=\tilde{t}_{I}$ in Eq. (66) is satisfactory for low pump energies, the increased thermal noise that is generated as the pump energy increases eventually forces one to alter the measurement times to account for the group velocity difference in the presence of this increased noise. To demonstrate this, we calculated the CV between the $n=3$ cavities in the signal and idler CROWS under a variety of pumping configurations and measurement times. In what follows, we fix $\tilde{t}_{S}$ and define the measurement delay time as $\tilde{\tau}=\tilde{t}_{S}-\tilde{t}_{I}$.
In Fig. 12, we plot the minimum and maximum CV as a function of $\kappa^{-1}$ for a fixed pump energy of 670 fJ for fourteen different $\tilde{\tau}$, from $\tilde{\tau}=0$ to $\tilde{\tau}=0.065\Gamma_{+}$ in increments of $0.005\Gamma_{+}$. Note that the minimum CV initially increases and then decreases as $\tilde{\tau}$ is increased, while the maximum CV monotonically decreases. The effect of varying $\kappa$ between $2\pi/15d$ and $2\pi/45d$ is less pronounced than that of varying the measurement time delay $\tilde{\tau}$. Specifically, with a fixed pump energy of 670 fJ (as in Fig. 12), we find that choosing $0.045\Gamma_{+}\leq\tilde{\tau}\leq 0.065\Gamma_{+}$ (the darkest six curves) greatly reduces the dependence of the CV minimum on $1/\kappa$, while also reducing the CV maximum. From an alternative perspective, if the optimum $\tilde{\tau}$ is chosen to minimize the CV minimum, then the frequency width (and thus the temporal width) of the pump pulse may be adjusted to minimize the CV maximum, which has little effect on the CV minimum.
In Fig. 13, we plot the minimum and maximum CV as a function of pump energy for a fixed $\kappa=2\pi/20d$. The pump energy has a more pronounced effect on the CV extrema than $\kappa$. Fig. 13 (b) shows that once the pump energy exceeds $\approx 1$ pJ, the entanglement vanishes unless the time-delay is sufficiently increased  [^2] The jogs in the CV minima in the lighter curves are caused by secondary peaks becoming more inseparable than the initial pulse for high enough pump energy.. From Fig. 13 (a), we see that for pulse energies above $\approx 1$ pJ, the CV maximum exhibits an exponential dependence on the pump energy: although entanglement may be recovered with an appropriate choice of $\tilde{\tau}$, the large increase in the CV maximum makes these higher pulse energies undesirable. We find that a good configuration that gives a low minimum CV, while keeping the pulse energy and the maximum CV modest, is as follows: a pump pulse energy of $\approx 670$ fJ, $\kappa=2\pi/20d$, and $\tilde{\tau}=0.45\Gamma_{+}$. This gives a minimum and maximum CV for the $n=3$ cavity pair of 0.923 and 1.74 respectively.
V Conclusion
We have presented a four-step theoretical approach to simulate the generation and propagation dynamics of a quantum state of light in lossy nanophotonic structures consisting of coupled open optical resonators. Using our theoretical approach, we designed a PC CCS for the generation of entangled 2MSTS of light on-chip and computed the time-dependent photon number and correlation variance in the electric field of this state at various points in space and time in the PC CCS. We used as a basis for our analysis the 184 leaky QMs of the full system, which were computed with a tight-binding approach. Our analysis showed that the RS, when excited with a Gaussian pump pulse carrying an initial energy between 0.5 and 1 pJ and a $k$-space width between $2\pi/15d$ and $2\pi/45d$, generates a two-mode squeezed thermal state, which exits the RS as two spatially separated pulses of light that remain entangled as they propagate.
We find that although the nonlinear generation process creates strong squeezing and low thermal photon noise in the RS, the squeezing degrades substantially after transmission into the output CROWs. The difference between the intrinsic quality factors of the RS QMs and the over-coupling of the pump CROW contribute to this degradation. Thus, it is clear that further refinement of the design of the RS and coupling to the CROWs in this system is required to balance the number of photons going into each of the CROWs, while keeping the loaded cavity Qs in the desired range.
Our analysis of the CV demonstrates the ability of our theoretical approach to efficiently and accurately calculate the generation and evolution of quantum states of light in coupled-cavity systems. The efficiency of our approach allows us to evaluate many pumping schemes and ascertain the effect of changes to the pump, or changes to any individual cavities in the system, at every stage of the process.
Although we focused on the time-dependent photon number and CV in this work, our method is general and can be applied to any observable which can be expanded in terms of the full-system QMs. The general procedure applied here may also be extended to any nanophotonic system composed of coupled elements whose modes are easily obtained. This approach is particularly useful when combining simple elements causes the entire system to lose the properties which allowed for efficient computation of the modes of the constituents.
Acknowledgements.

### v3

@@ Appendix D Quantum State Evolution in the Output CROWs
In this Appendix, we model the free quantum evolution of the 2MSTS in the full system in the presence of loss but without the pump and compute the time-dependent expectation value of the CV between output CROW cavities. We accomplish this by first solving the adjoint master equation for the time-evolution of the full-system QM ladder operators, and then use the initial state of the system at time $\tilde{t}=\tilde{t}_{0}$ to calculate time dependence of the photon number and correlation variance.
@@
Consider the CV, $\Delta^{2}(p_{S},\tilde{t}_{S};p_{I},\tilde{t}_{I})$ between two cavities, one each in the signal and idler CROW, denoted $p_{S}$ and $p_{I}$ respectively (see Fig. 1), where we allow the measurement times $\tilde{t}_{S}$ and $\tilde{t}_{I}$ to be different. This CV may be expressed in terms of the quadrature operators in the two cavities as
$\Delta^{2}(p_{S},\tilde{t}_{S};p_{I},\tilde{t}_{I})=$ (72)
$\expectationvalue{\quantity[X_{p_{S}}(\tilde{t}_{S})-X_{p_{I}}(\tilde{t}_{I})]^{2}}$
$+\expectationvalue{\quantity[Y_{p_{S}}(\tilde{t}_{S})+Y_{p_{I}}(\tilde{t}_{I})]^{2}},$
@@
Below, we shall write the time-dependent photon number and correlation variance in terms of the $a$-operators. First, though, we must determine the time-evolution of the $c$-operators, since, as we shall see, it is these (and not the $b$-operators) that display simple single-frequency time-dependence. We begin with the adjoint master equation, which describes the time-evolution of a Heisenberg picture operator $A_{H}$ in terms of the free Hamiltonian in Eq. (22) [KamandarDezfouli2014]:
@@
[…]
$=\stackrel{{\scriptstyle\leftrightarrow}}{{\Phi}}(\tilde{t})\vec{a},$ (79)
where $\vec{a}\equiv\quantity(a_{1},a_{2},...,a_{p})^{T}$ and $\stackrel{{\scriptstyle\leftrightarrow}}{{F}}(\tilde{t})=\text{diag}\quantity(e^{-i\tilde{\omega}_{1}(\tilde{t}-\tilde{t}_{0})},e^{-i\tilde{\omega}_{2}(\tilde{t}-\tilde{t}_{0})},...,e^{-i\tilde{\omega}_{\nu}(\tilde{t}-\tilde{t}_{0})})$. Inserting Eq. (34) into Eqs. (73, 74) and substituting the result into Eq. (72) yields, after some algebra
$\Delta^{2}(p_{S},\tilde{t}_{S};p_{I},\tilde{t}_{I})=$ (80)
$\frac{1}{2}(B_{p_{S}p_{S}}+B_{p_{I}p_{I}})$
$+\sum_{q,q^{\prime}=1}^{N}[\phi_{p_{S}q}^{*}(\tilde{t}_{S})\phi_{p_{S}q^{\prime}}(\tilde{t}_{S})$
$+\phi_{p_{I}q}^{*}(\tilde{t}_{I})\phi_{p_{I}q^{\prime}}(\tilde{t}_{I})]\expectationvalue{a^{\dagger}_{q}a_{q^{\prime}}}$
[…]
where $p$ and $q$ are now cavities in the RS, the $d_{j}$ ($d_{j}^{\dagger}$) are the annihilation (creation) operators for the $j^{\text{th}}$ RS QM, and the $\sigma_{pj}$ are the matrix elements of the matrix $\stackrel{{\scriptstyle\leftrightarrow}}{{\Sigma}}=\stackrel{{\scriptstyle\leftrightarrow}}{{U}}^{-1}$ (see Eq. (54). In Eqs. (81-82), $j$ and $k$ run through only the signal ($j=S$) and idler ($k=I$) modes of the RS.
We now directly evaluate the ensemble averages in Eqs. (81, 82). Using the density operator from Eq. (10), the ensemble average for a general operator $A$, acting on a two-mode squeezed state can be expressed as:
[…]
Using the standard relations [GerryKnight], we obtain:
@@
Thus, from Eq. (83) and Eqs. (84 - 87), we can evaluate Eqs. (81, 82). After some algebra, we obtain:
$\expectationvalue{a_{p}a_{q}}=$ (88)
$-\quantity({\sigma}^{*}_{pS}{\sigma}^{*}_{qI}+{\sigma}^{*}_{pI}{\sigma}^{*}_{qS})$
$\times\quantity(n^{th}_{S}+n^{th}_{I}+1)e^{i\theta}\cosh r,\sinh r,$
and
$\expectationvalue{a_{p}^{\dagger}a_{q}}=$ (89)
$\sigma_{pS}{\sigma}^{*}_{qS}\quantity(n^{th}_{S}\cosh^{2}r+(1+n^{th}_{I})\sinh^{2}r)$
$+\sigma_{pI}{\sigma}^{*}_{qI}\quantity(n^{th}_{I}\cosh^{2}r+(1+n^{th}_{S})\sinh^{2}r),$
@@
where $r$, $n^{\text{th}}_{S}$, and $n^{\text{th}}_{I}$ are evaluated at $\tilde{t}=\tilde{t}_{0}$. Finally, the combination of Eqs. (88, 89) with Eq. (80) yields an expression for the correlation variance between two cavities in the signal and idler output CROWs. The preceding analysis also yields an expression for the average photon number in the output CROW cavities. Since Eq. (38) is normally ordered, we may simply use Eq. (79) to obtain Eq. (36).